Four persons can hit a target correctly with probabilities and respectively. If all hit at the target independently, then the probability that the target would be hit, is
- A.
- B.
- C.
- D.
Show answer and solution
Correct option: B
Solution
P(target is hit) = 1 - P(none hits) = 1 - (1 - 1/2)(1 - 1/3)(1 - 1/4)(1 - 1/8) = 1 - (1/2)(2/3)(3/4)(7/8) = 1 - 42/192 = 1 - 7/32 = 25/32.
[JEE Main 2019 · 9 Apr · Shift II]